OpenStudy (anonymous):

√16 - 8 in simplified radical form?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \sqrt{16}-8~, }\)\(\LARGE\color{blue}{ \bf ~~or~~is~~it~~\LARGE\color{blue}{ \bf \sqrt{16-8} }}\)

OpenStudy (anonymous):

the second one! ah i didnt even know there was a difference, i apologize

OpenStudy (anonymous):

Subtract, then divide/simplify. \[2\sqrt{2}\]

OpenStudy (anonymous):

√(16-8) signifies what you're asking.

OpenStudy (whpalmer4):

\[\sqrt{16-8} = \sqrt{8} = \sqrt{2*2*2} = \sqrt{2*2} * \sqrt{2} = 2*\sqrt{2}=2\sqrt{2}\] For every pair of identical factors under the square root sign, you can replace the pair with 1 outside of the square root sign, as I did here with 2*2